OpenStudy (anonymous):

The area of a rectangle is given by A = x2 + 4x – 5. Find the value of x if the total area of the rectangle is equal to 7 square units. a) x = 1 b) x = 2 c) x = 7 d) x = 5

OpenStudy (anonymous):

Set A=7 and solve for x. \[x^{2}+4x-5=7\]\[x^{2}+4x-12=0\]\[(x+6)(x-2)=0\] Only one answer will make sense.